OpenStudy (anonymous):

A bag contains 7 red marbles, 12 orange marbles, 18 yellow marbles, and 5 purple marbles. A marble is drawn at random from the bag. What is the probability that the marble drawn will not be orange?

OpenStudy (anonymous):

add all marbles together 7+12=19+5=24+18=42 now subtract the number of orange marbles 42-12=30 then make ratio out of it 30/42

OpenStudy (anonymous):

thank you :)